East Coast Counties Midwest Counties The table to the right gives the cost of living index (COLI) for six East Coast countiesc. Find the standard deviation for each of the two data sets. The standard deviation for the East Coast Counties is : (Type aHow well does the rule work in each case? Briefly discuss why it does or does not work well. Choose the correct answer below.
Answer #1

C. East : sd =Li=1 EX (.; - 72 N-1 = 77.7836 Midwest: sd - Liziri - 02 N 1 = 55.4197 d. East : sd = range/4= 210/4 = 52.5 MW:

Dear student, if you find this useful, give your feedback through voting.